what is the diameter of a 1.00m lenght of tungsten wire whose resistance is 0.18 ohm?.2 significant figures
Answer #1

l=1.0m R = 0.182 8 = 5.60 8 10 8 r.m. using R = SA Aase - Tr? - et d=28 = d=2 / se VRTT d=2 / 5.60x100x10 0.18 * 3.14 d = 6:2